OpenStudy (anonymous):

equation of a line perpendicular to y=-1/3x+1?. passing through the point (2,7)

OpenStudy (anonymous):

The linear relationship between two perpendicular lines is that $m_1 = -\frac 1 m$. In other words, the second slope will be the negative multiplicative inverse of the first one.

OpenStudy (anonymous):

Gah. \(m_1 = -\frac 1 m\).

OpenStudy (anonymous):

Thus, the slope is 3. Now, we have to find $b$. $$y = mx + b \\ y = 3x + b 7 = 3(2) + b 7 = 6 + b b = 1$$ Thus, the equation is \(y = 3x + 1\).

OpenStudy (anonymous):

Sorry for the bad formatting. I'm trying to get adapted to not using math.stackexchange.com.

OpenStudy (anonymous):

No problem! But if the given solutions are A. 3x - y = -1 B. -3x + y = -1 C. 3x - y = 1 D. -3x - y = 1 Then would the answer be B, or C? Or maybe something else entirely? @bloopman

OpenStudy (anonymous):

Ah, well, now we have to put it in standard form. $$y = 3x + 1 \\ -3x + y = 1$$ The answer is therefore B.

OpenStudy (anonymous):

Note all I did was subtract $3x$ from the right side.